# Germany, https://ais.uni-bonn.de

### v1

@@ 1 Introduction
With the rapid evolution of deep learning research, particularly in the natural language domain, we have witnessed the emergence of enormous Transformer-based models capable of generating high-quality texts and synthetic images [brown2020language, ramesh2022hierarchical]. They hold much potential for open-ended robotic applications, like domestic service tasks, but cannot be directly applied, because robotics involves additional modalities, such as images and actions.
In 2021, OpenAI released the foundation model CLIP, which uses a contrastive learning paradigm to connect text and images [radford2021learning]. Representations of textual and visual inputs are learned by contrasting similar and dissimilar pairs. CLIP features are suitable for open-vocabulary zero-shot image categorization.
Many robotics applications soon utilized CLIP-based models for different open-vocabulary tasks [shafiullahclip]. Vision-Language-Action models like OpenVLA [OpenVLA] predict the next action based on visual inputs. Research in models predicting entire action transcript sequences based on language instructions and visual inputs is still in its early phase. Training a model that understands all relevant modalities and has strong generalizability is challenging and requires a considerable amount of data.
In this work, we introduce the end-to-end model LIAM, which receives inputs from the modalities Language, Images, Actions, and Maps and predicts action sequences. Two CLIP-like self-supervised training paradigms are designed to pre-align two or more different embedding spaces during the pre-training stage, enhancing the model’s ability to understand and process multimodal inputs. We utilized semantic maps as an additional modality directly in addition to textual and visual input, where the map modality is designed to serve as a knowledge base for the model. We evaluate our model with the ALFRED Challenge [shridhar2020alfred], a benchmark task that requires a robot to understand and execute complex natural language instructions in a simulated environment. The code and pre-trained models to reproduce the results are made publicly available: https://github.com/AIS-Bonn/LIAM.
@@ 2 Related Work
Open-vocabulary approaches have garnered significant attention in the field of computer vision. In the past decade, most vision downstream tasks were trained via datasets with a fixed number of classes. The CLIP model, with its contrastive learning paradigm, brings a promising solution for inferencing open vocabulary vision Tasks. After embedding the text and image using separate encoders, CLIP is trained by maximizing the similarity score of the representation of corresponding image and text pairs [radford2021learning].
A massive amount of work is built on CLIP models for different down-streaming tasks, e.g., object detection [zhou2022detecting, gu2021open], semantic segmentation [luddecke2022image, xu2022groupvit], action prediction [wang2021actionclip, luo2022clip4clip]. These approaches have designed similar methods adapted to their specific vision downstream tasks. For example, GroupViT first uses a group-based approach for semantic segmentation by introducing learnable group tokens to the ViT architecture ([xu2022groupvit]). During inference time, each group is assigned the semantic label with the highest similarity score representing the text input.
In order to tackle the Vision-Language Navigation (VLN) and Vision-Language Manipulation (VLM) tasks, there are two main approaches. I. An End-to-End model trains a model that can accept all modal inputs simultaneously and make predictions of current action for the agent [shridhar2020alfred, pashevich2021episodic]. For example, Episodic Transformer (E.T.) is an end-to-end attention-based transformer model that consists of three modality-specific encoders and one multi-modality fusion encoder [pashevich2021episodic]. II. Modular Methods propose a pipeline of learned modules for tackling the challenge of understanding environment and the tasks [blukis2022persistent, min2021film]. FILM proposes a language processing module, which converts the task description into an action transcript using pre-defined templates. A semantic mapping module processes ego-centric RGB images into a 2d semantic map. Furthermore, a semantic search policy predicts the possible distribution of objects’ locations. A deterministic policy is designed to predict the action decision based on the predicted action transcript and the predicted semantic map.
@@ 3 Main Approach
We now introduce the LIAM model, including the pre-training stage and the action generation. An overview of the model is given in Figure 1. The model architecture is adopted from the Episodic Transformer [pashevich2021episodic], but replace the image and text encoders with CLIP encoders and employ a semantic map [min2021film]. To manage the high computational cost, we freeze the weights and biases of the CLIP backbone during end-to-end model training, ensuring efficient resource utilization.
All experiments are conducted using ALFRED (Action Learning From Realistic Environments and Directives) [shridhar2020alfred], a recent simulator generated benchmark for learning a mapping task, from vision (ego-centric RGB image) and natural language input to an action transcript for a domestic service robot. ALFRED’s challenge provides the following annotations: an initial state of the simulated room, language instructions, and an expert demonstration trajectory. Language instructions are provided as global instructions, e.g., "Pick up the alarm clock and turn on the lamp. <<much-less-than$<$ <goal>>much-greater-than$>$ >" and a list of sentences for sub-goal instructions. The annotation consists of a sequence of discrete actions, the object’s mask, whenever the interaction with objects is involved.
@@ 3.1 Contrastive Alignment Pre-training
CLIP was initially trained using data gathered from the Internet, resulting in bias and significant performance differences across different datasets in zero-shot inference [radford2021learning]. Given that the ALFRED dataset is simulator-generated, it deviates slightly from reality. Hence, a pre-training stage to align the latent space of image and action becomes imperative.
@@
In our approach, we used f(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ), a 1D convolution layer along with global average pooling, to compute a fused representation of two consecutive images. We embedded actions with an embedding layer that embeds all 14 classes (12 classes + 2 special tokens) into 768-dimensional dense vectors. After having the normalized representations of both frames $I_{e}$ and action classes $A_{e}$, we compute a cosine similarity score $s$ of both vectors and fine-tuned the CLIP backbone following CLIP’s approach. With $a_{i}\in A_{e},I_{t}\in I_{e}$, the similarity matrix P is defined as:
@@
We notated text, action, and image with alphabet T, A, and I. I2A denotes "image-to-action," and all later mentioned abbreviations (X2X) follow this pattern. Function f(,)($\cdot$,$\cdot$)( ⋅ , ⋅ ) denotes the fusion function of two consecutive images at time step t and t + 1 as described above. $\tau$ is the temperature, a learnable parameter scales the similarity score. Following CLIP, we initialized the temperature with a value of 0.07 and use 100 as an upper threshold to clip the temperature value during the training stage, preventing the scaling factor from becoming too large [wang2021actionclip] and 0.01 as a lower threshold bound to prevent dividing by zero error.
Following ActionCLIP [wang2021actionclip], we used the KL divergence loss to fine-tune the CLIP backbone. The total loss of the contrastive alignment pre-training is the average of both image-to-action and action-to-image loss:
@@ 3.2 Triple Contrastive Pre-training
Our contrastive learning pre-aligns the latent space of the vision and action embedding. However, the alignment of the language embedding space with other latent spaces is thus neglected. To address this, we introduced a triple contrastive pre-training stage. This method not only aligns the vision and action embedding spaces but also aligns the language embedding space with them. This mutual training ensures that the loss can be back-propagated to all encoders. In addition to Figure 3, Figure 3 shows the example of the ground truth of aligning the CLIP text encoder and image encoder. The training objective is straightforward: the representation of the language instruction and its corresponding frame sequence should be similar to each other. We first computed the representation of every single frame (N frames) using the CLIP image encoder and further process this sequence of representations in two ways.
@@
where S is the similarity matrix.
The total loss of triple contrastive alignment is defined as:
@@ 3.3 Action Generation
We embedded text and visual input using independent CLIP encoders, which are now pre-aligned. The map is embedded using a trainable projection layer (FC layer + GELU). We first applied positional encoding and modal-type encoding to each representation. The learnable modal-type encoding, first introduced by VILT [kim2021vilt], brings extra information and clarifies different modality types to the Transformer model; it also avoids the need for explicit separation tokens like “[SEP].”
After all the representations is concatenated, we fed it further to a multi-modality fusing layer (2-layer Transformer encoder) to learn a general representation. The multimodal Transformer is a comprehensive model that aims to integrate all different modalities into one global representation, namely language, visual frames, action sequence, and semantic map in our case. We use causal attention following E.T. [pashevich2021episodic]. The attention mask follows these rules: Language tokens should attend to the language itself. Visual frames are allowed to attend to all language tokens, but they can only attend to all the frames, actions, and semantic maps before current time step t. The same rules are applied to action tokens and semantic maps. Sinusoidal positional encoding is also applied.
@@ 3.4 LIAM Pipeline
In this section, we define our input and output mathematically to explain our approach more deeply. Given the following inputs: Tokenized language instruction $[L_{1},…L_{m}]$, Tokenized language instruction $[I_{1},...,I_{n}]$, Action sequence $[A_{1},…,A_{n}]$, Semantic map sequence $[M_{1},…,M_{n}]$, where language tokens have length m, and all other modalities have length n. For an image sequence $I_{1},\dots,I_{n}$, only n - 1 actions are involved. We append "<<stop>>" as the last action. Four different encoders encode all input to their latent space and output $F_{l},F_{v}$,$F_{a}$$\&F_{m}$ as the feature representation.
[…]
@@
$O=\overline{F}^{d}_{I}D$
After applying modal-type and positional encoding to each feature representation, we concatenate them into $F_{total}^{0}$. The multi-modal Transformer encoder has $\mathbb{L}$ layers, each with Multi-headed Attention (MA), Layer Normalization (LN), and residual connections. We then extract the visual part of the total representation (from index m + 1 to m + 1 + n) for action sequence prediction. The final action sequence is obtained by multiplying this sliced matrix with a trained linear layer D.
The end-to-end model is trained using cross-entropy loss between the generated action sequence and the ground truth action sequence. For predicted action sequence P = [$\hat{a_{1}}\dots\hat{a_{n}}$] and ground truth Q = [$a_{1}\dots a_{n}$], the cross-entropy loss is defined as:
$\mathcal{L}_{\text{action}}=-\sum^{n}_{i=1}\sum^{13}_{c=1}a_{i,c}log(\hat{a_{i,c}}),$
@@
Following E.T. [pashevich2021episodic], we also use auxiliary losses in the training. First, we predict the object class for each frame ("NoObject" is also a class). The predicted object class is a list with the same length of the action. Similary to the action, for predicted object sequence $P_{o}$ = [$\hat{o_{1}}\dots\hat{o_{n}}$] and ground truth Q = [$o_{1}\dots o_{n}$], the cross-entropy loss is defined as:
$\mathcal{L}_{\text{object}}=-\sum^{n}_{i=1}\sum^{85}_{c=1}o_{i,c}log(\hat{o_{i,c}}),$
@@
$\mathcal{L}_{total}=\mathcal{L}_{action}+\alpha\mathcal{L}_{object}+\beta\mathcal{L}_{gp}.$
@@ 4.1 Pre-training Stage
Table 1 presents the accuracy score of correct matching pairs among the mini-batches. Firstly, we can clearly see that no CLIP model has the ability to correctly match two consecutive visual images with their corresponding actions in a zero-shot manner. As discussed before, CLIP is a model that is sensitive to the dataset. However, both image-action contrastive and triple contrastive alignment can be adequately trained and show significant accuracy gains in matching corresponding pairs of two consecutive images and action or pairs of the image sequence and text sequence. We use batch size 64 for image-action contrastive alignment training, batch size three, and additionally, constraint each sequence to length 21 (during validation) for triple-contrastive alignment to keep a fair comparison between both accuracy scores. We observed that using zero-shot inference, the accuracy for text sequence and image sequence matching is around 33%, and the accuracy for two image frames and action matching lies between 3% and 8%. Both scores denote the random guessing from the model. After fine-tuning, the accuracy of predicting the correct image and text sequence pair is increased using MobileCLIP backbone and CLIP VIT-B-32 backbone to 60.22% and 45.62%, however still not entirely correct out of the mini-batch. We observe that triple-contrastive alignment indeed benefits accuracy in matching the correct action label of the consecutive images using all three backbones. Overall, based on its performance in the action-image matching task, we decide to use pre-aligned CLIP VIT-B-32 as the backbone to extract the features of our input.
@@ 4.2 Quantitative Results
Table 2: Results for end-to-end model predicted sequence and the ground truth.
@@
In our experiments, we used OpenCLIP with ViT B-32 as our backbone, which we found to be the most suitable in our experiments. From Table 2, first, we observe that the fine-tuning of the CLIP backbone brought benefits in increasing the accuracy and the F1 score. Although the improvement in accuracy is not significant, e.g., 73.00%, 76.97%, 77.52% in the seen data, 71.87%, 79.56%, 76.89% in the unseen data, respectively. The improvement in F1 score is relatively huge, 8% in the seen data, and around 15% improvement in unseen data, when the triple contrastive pre-alignment is performed. Second, the semantic map improved the model’s performance for the zero-shot backbone and triple-aligned backbone, showing the potential of using the map as an extra modality to enhance the model’s spatial understanding rather than solely given the ego-centric RGB images. The backbone fine-tuned with image-to-action alignment has an acceptable accuracy score but a relatively low F1 score; this indicates that the model still suffers from the imbalanced data problem.
@@ 4.3 Qualitative Results
Figure 4 shows the decision of our agent for the task "Pick up the basketball and turn on the desk lamp in the bedroom." The agent can comprehend most of the sub-goal instructions and correctly find and pick up the basketball at the foot of the bed. Furthermore, the agent rotates left twice and goes to the desk. However, this task eventually failed because of the model’s wrong prediction of the lamp’s location (which should be on the other side of the desk).

### v2

# LIAM: Multimodal Transformer for Language Instructions, Images, Actions and Semantic Maps

@@ 1 Introduction
With the rapid evolution of deep learning research, particularly in the natural language domain, we have witnessed the emergence of enormous Transformer-based models capable of generating high-quality texts and synthetic images [1, 2]. They hold much potential for open-ended robotic applications, like domestic service tasks, but cannot be directly applied because robotics involves additional modalities, such as images and actions.
In 2021, OpenAI released the foundation model CLIP, which uses a contrastive learning paradigm to connect text and images [3]. Representations of textual and visual inputs are learned by contrasting similar and dissimilar pairs. CLIP features are suitable for open-vocabulary zero-shot image categorization.
Many robotics applications soon utilized CLIP-based models for different open-vocabulary tasks [4]. Vision-Language-Action models like OpenVLA [5] predict the next action based on visual inputs. Research in models predicting entire action transcript sequences based on language instructions and visual inputs is still in its early phase. Training a model that understands all relevant modalities and has strong generalizability is challenging and requires a considerable amount of data.
In this work, we introduce the end-to-end model LIAM, which receives inputs from the modalities Language, Images, Actions, and Maps and predicts action sequences. Two CLIP-like self-supervised training paradigms are designed to pre-align two or more different embedding spaces during the pre-training stage, enhancing the model’s ability to understand and process multimodal inputs. We utilized semantic maps as an additional modality directly in addition to textual and visual input, where the map modality is designed to serve as a knowledge base for the model. We evaluate our model with the ALFRED Challenge [6], a benchmark task that requires a robot to understand and execute complex natural language instructions in a simulated environment. The code and pre-trained models to reproduce the results are made publicly available: https://github.com/AIS-Bonn/LIAM.
@@ 2 Related Work
Open-vocabulary approaches have garnered significant attention in the field of computer vision. In the past decade, most vision downstream tasks were trained via datasets with a fixed number of classes. The CLIP model, with its contrastive learning paradigm, brings a promising solution for inferencing open vocabulary vision tasks. After embedding the text and image using separate encoders, CLIP is trained by maximizing the similarity score of the representation of corresponding image and text pairs [3].
A massive amount of work is built on CLIP models for different down-streaming tasks, e.g., object detection [7, 8], semantic segmentation [9, 10], action prediction [11, 12]. These approaches have designed similar methods adapted to their specific vision downstream tasks. For example, GroupViT first uses a group-based approach for semantic segmentation by introducing learnable group tokens to the ViT architecture ([10]). During inference time, each group is assigned the semantic label with the highest similarity score representing the text input.
To tackle the Vision-Language Navigation (VLN) and Vision-Language Manipulation (VLM) tasks, there are two main approaches. I. An End-to-End model trains a model that can accept all modal inputs simultaneously and make predictions of current action for the agent [6, 13]. For example, Episodic Transformer (E.T.) is an end-to-end attention-based transformer model that consists of three modality-specific encoders and one multi-modality fusion encoder [13]. II. Modular Methods propose a pipeline of learned modules for tackling the challenge of understanding the environment and the tasks [14, 15]. FILM proposes a language processing module, which converts the task description into an action transcript using pre-defined templates. A semantic mapping module processes egocentric RGB images into a 2D semantic map. Furthermore, a semantic search policy predicts the possible distribution of objects’ locations. A deterministic policy is designed to predict the action decision based on the predicted action transcript and the predicted semantic map.
@@ 3 Main Approach
We now introduce the LIAM model, including the pre-training stage and the action generation. An overview of the model is given in Figure 1. The model architecture is adopted from the Episodic Transformer [13], but we replace the image and text encoders with CLIP encoders and employ a semantic map [15]. To manage the high computational cost, we freeze the weights and biases of the CLIP backbone during end-to-end model training, ensuring efficient resource utilization.
All experiments are conducted using ALFRED (Action Learning From Realistic Environments and Directives) [6], a recent simulator generated benchmark for learning a mapping task, from vision (egocentric RGB image) and natural language input to an action transcript for a domestic service robot. ALFRED’s challenge provides the following annotations: an initial state of the simulated room, language instructions, and an expert demonstration trajectory. Language instructions are provided as global instructions, e.g., "Pick up the alarm clock and turn on the lamp. <<<<goal>>>>" and a list of sentences for sub-goal instructions. The annotation consists of a sequence of discrete actions, the object’s mask, whenever the interaction with objects is involved.
@@ 3.1 Contrastive Alignment Pre-training
CLIP was initially trained using data gathered from the Internet, resulting in bias and significant performance differences across different datasets in zero-shot inference [3]. Given that the ALFRED dataset is simulator-generated, it deviates slightly from reality. Hence, a pre-training stage to align the latent space of image and action becomes imperative.
@@
In our approach, we used f(,)($\cdot$$\cdot$), a 1D convolution layer along with global average pooling, to compute a fused representation of two consecutive images. We embedded actions with an embedding layer that embeds all 14 classes (12 classes + 2 special tokens) into 768-dimensional dense vectors. After having the normalized representations of both frames $I_{e}$ and action classes $A_{e}$, we computed a cosine similarity score $s$ of both vectors and fine-tuned the CLIP backbone following CLIP’s approach. With $a_{i}\in A_{e},I_{t}\in I_{e}$, the similarity matrix P is defined as:
@@
We notated text, action, and image with the letters T, A, and I. I2A denotes "image-to-action," and all later mentioned abbreviations (X2X) follow this pattern. Function f(,)($\cdot$$\cdot$) denotes the fusion function of two consecutive images at time step t and t + 1 as described above. $\tau$ is the temperature, a learnable parameter scales the similarity score. Following CLIP, we initialized the temperature with a value of 0.07 and use 100 as an upper threshold to clip the temperature value during the training stage, preventing the scaling factor from becoming too large [11] and 0.01 as a lower threshold bound to prevent dividing by zero error.
Following ActionCLIP [11], we used the KL divergence loss to fine-tune the CLIP backbone. The total loss of the contrastive alignment pre-training is the average of both image-to-action and action-to-image loss:
@@ 3.2 Triple Contrastive Pre-training
Our contrastive learning pre-aligns the latent space of the vision and action embedding. However, the alignment of the language embedding space with other latent spaces is neglected. To address this, we introduced a triple contrastive pre-training stage. This method not only aligns the vision and action embedding spaces but also aligns the language embedding space with them. This mutual training ensures that the loss can be back-propagated to all encoders. In addition to Figure 3, Figure 3 shows the example of the ground truth of aligning the CLIP text encoder and image encoder. The training objective is straightforward: the representation of the language instruction and its corresponding frame sequence should be similar to each other. We first computed the representation of every single frame (N frames) using the CLIP image encoder and further processed this sequence of representations in two ways.
@@
where S is the similarity matrix. The total loss of triple contrastive alignment is defined as:
@@ 3.3 Action Generation
We embedded text and visual input using independent CLIP encoders, which are now pre-aligned. The map is embedded using a trainable projection layer (FC layer + GELU). We first applied positional encoding and modal-type encoding to each representation. The learnable modal-type encoding, first introduced by VILT [16], brings extra information and clarifies different modality types to the Transformer model; it also avoids the need for explicit separation tokens like “[SEP].”
After all the representations are concatenated, we fed it further to a multi-modality fusing layer (2-layer Transformer encoder) to learn a general representation. The multimodal Transformer is a comprehensive model that aims to integrate all different modalities into one global representation, namely language, visual frames, action sequence, and semantic map in our case. We use causal attention following E.T. [13]. The attention mask follows these rules: Language tokens should attend to the language itself. Visual frames are allowed to attend to all language tokens, but they can only attend to all the frames, actions, and semantic maps before the current time step t. The same rules are applied to action tokens and semantic maps. Sinusoidal positional encoding is also applied.
@@ 3.4 LIAM Pipeline
In this section, we define our input and output mathematically to explain our approach more deeply. Given the following inputs: Tokenized language instructions $L_{1},…L_{m}]$, image sequence $I_{1},...,I_{n}]$, action sequence $A_{1},…,A_{n}]$, semantic map sequence $M_{1},…,M_{n}]$, where language tokens have length m, and all other modalities have length n. For an image sequence $I_{1},\dots,I_{n}$, only n - 1 actions are involved. We append "<<stop>>" as the last action. Four different encoders encode all input to their latent space and output $F_{l},F_{v}$$F_{a}$ and $F_{m}$ as the feature representation:
$\overline{F}_{mod}=F_{mod}+mod^{Type}+mod^{Pos},\hskip 5.69054pt\text{for mod in L, I, A, M.}$
@@
$O=\overline{F}^{d}_{I}D.$
After applying modal-type and positional encoding to each feature representation, we concatenate them into $F_{total}^{0}$. The multi-modal Transformer encoder has $\mathbb{L}$ layers, each with Multi-headed Attention (MA), Layer Normalization (LN), and residual connections. We then extract the visual part of the total representation (from index m + 1 to m + 1 + n) for action sequence prediction. The final action sequence is obtained by multiplying this sliced matrix with a trained linear layer D. The end-to-end model is trained using cross-entropy loss between the generated action sequence and the ground truth action sequence. For predicted action sequence P = $\hat{a_{1}}\dots\hat{a_{n}}$] and ground truth Q = $a_{1}\dots a_{n}$], the cross-entropy loss is defined as:
[…]
@@
Following E.T. [13], we also use auxiliary losses in the training. First, we predict the object class for each frame ("NoObject" is also a class). The predicted object class is a list with the same length as the action. Similary to the action, for predicted object sequence $P_{o}$ = $\hat{o_{1}}\dots\hat{o_{n}}$] and ground truth Q = $o_{1}\dots o_{n}$], the cross-entropy loss is defined as:
[…]
@@
$\mathcal{L}_{total}=\mathcal{L}_{action}+\alpha\mathcal{L}_{object}+\beta\mathcal{L}_{gp},$
@@ 4.1 Pre-training Stage
Table 1 presents the accuracy score of correct matching pairs among the mini-batches. Firstly, we can clearly see that no CLIP model can correctly match two consecutive visual images with their corresponding actions in a zero-shot manner. As discussed before, CLIP is a model that is sensitive to the dataset. However, both image-action contrastive and triple contrastive alignment can be adequately trained and show significant accuracy gains in matching corresponding pairs of two consecutive images and action or pairs of the image sequence and text sequence. We use batch size 64 for image-action contrastive alignment training, batch size three, and additionally, constraint each sequence to length 21 (during validation) for triple-contrastive alignment to keep a fair comparison between both accuracy scores. We observe that using zero-shot inference, the accuracy for text sequence and image sequence matching is around 33%, and the accuracy for two image frames and action matching lies between 3% and 8%. Both scores denote the random guessing from the model. After fine-tuning, the accuracy of predicting the correct image and text sequence pair is increased using MobileCLIP backbone and CLIP VIT-B-32 backbone to 60.22% and 45.62%, however still not entirely correct out of the mini-batch. We observe that triple-contrastive alignment indeed benefits accuracy in matching the correct action label of the consecutive images using all three backbones. Overall, based on its performance in the action-image matching task, we decided to use pre-aligned CLIP VIT-B-32 as the backbone to extract the features of our input.
@@ 4.2 Quantitative Results
Table 2: Results for the end-to-end model predicted sequence and the ground truth.
@@
In our experiments, we used OpenCLIP with ViT B-32 as our backbone, which we found to be the most suitable in our experiments. From Table 2, first, we observe that the fine-tuning of the CLIP backbone brought benefits in increasing the accuracy and the F1 score. Although the improvement in accuracy is not significant, e.g., 73.00%, 76.97%, 77.52% in the seen data, 71.87%, 79.56%, 76.89% in the unseen data, respectively. When the triple contrastive pre-alignment is applied, the F1 score improves by 8% on seen data and by around 20% on unseen data, compared to the contrastive alignment approach. Second, the semantic map improved the model’s performance for the zero-shot backbone and triple-aligned backbone, showing the potential of using the map as an extra modality to enhance the model’s spatial understanding rather than solely given the egocentric RGB images. The backbone fine-tuned with image-to-action alignment has an acceptable accuracy score but a relatively low F1 score; this indicates that the model still suffers from the imbalanced data problem.
@@ 4.3 Qualitative Results
Figure 4 shows the decision of our agent for the task, "Pick up the basketball and turn on the desk lamp in the bedroom." The agent can comprehend most of the sub-goal instructions and correctly find and pick up the basketball beneath the bed. Furthermore, the agent rotates left twice and goes to the desk. However, this task eventually failed because of the model’s wrong prediction of the lamp’s location (which should be on the other side of the desk).